2a [5 marks] Click on the link and open the pdf. (if you zoom in, you can see the gridlines). If the market price is $4.50, calculate the firm’s economic profits or losses in the short-run.

2b [5 marks] You learn that in the market at this price the total quantity transacted is equal to 7,040. How many firms are operating in the market? What are the industry profits/losses?

2c [5 marks] If the firm in the graph is representative for all other firms in the industry, what will be the market price in the long run? Describe briefly how the market moves from the short-run to the long-run.

ATC AVC OTTTT 0 1 2 3 4 5 6 7
Answer #1

a) P=4.5

Setting P=MC, the firm will produce Q = 5.5 units where ATC = 3

So, Profits = P-ATC*Q = (4.5-3)*5.5 = 8.25

b) A sigle firm supplies 5.5 units so total number of firms = 7040/5.5 = 1280

Industry profit = (4.5-3)*7040 = 10560

c) Market price in the long run is equal to the minimum ATC = 3 where P=ATC

In the short run, since the industry is making positive economic profits, this will attract more firms to the market and market supply will increase.When supply increases, the price decreases until it is equal to the minimum ATC where all the firms are earning zero profits so that no new firms will have any incentive to enter the market.